A total of $21,000 is to be​ invested, some in bonds and some in certificates of deposit​ (CDs). If the amount invested in bonds is to exceed that in CDs by $3,000, how much will be invested in each type of​ investment?

- this is what I am given I need to find the amount invested is CDs and amount invested in bonds. my book says something about certification interest but I'm confused on how I should set up the problem bc I got 9000. do I set it up (b+3000)=21000? what do I do? plz help

Answer by ikleyn(52943)

.
Let x = the amount invested in bonds and y = the amount invested in CDs.


x + y = 21000     (1)

x - y =  3000     (2)


Add the equations to get


2x = 21000 + 3000 = 24000  ====>  x = 24000/2 = 12000.


Then from equation (1),  y = 21000 - 12000 = 9000.


ANSWER.  $12000 in bonds  and  $9000 in CDs.

Solved.